3) The presence of spots in a ladybug species is controlled by a single locus. The allele S produces spots while the allele s produces no spots. A breeding study shows that S is completely dominant to s. In a random sample of 758 individuals, 28 completely lack spots.

a) What are the expected frequencies of the three genotypes in this population and what assumption(s) did you make to solve this problem?

b) What are the allele frequencies?

4) In some species of goats, the presence of horns is controlled by a single locus and determined by an allele that is dominant in males and recessive in females. It is not sex-linked. If 96% of males have horns, what proportion of females have horns. Use the alleles H (horns) and h (no horns). Show your work.
Answer #1

3.If you assume the population is in HWC equilibrium then, q2 = 28/758 = 0.037, q = (0.037)1/2 = 0.192, p = 1 – q = 0.808 Expected HWC genotype frequencies p2 = (0.808)2 = 0.653 2pq = 2(0.808)(0.192) = 0.311 q2 = (0.192)2 = 0.037